OpenStudy (anonymous):

differentation find increasing,decreasing,maximum and minimum f(x)=x-13/x^2-25

OpenStudy (anonymous):

\[f(X)=\frac{ x-13 }{ x^{2}-25 }\]

OpenStudy (anonymous):

I guide you only, take derivative first, tell me the result

OpenStudy (anonymous):

1. differentiate using quotient rule OR 2. convert the fraction to a product and use product rule.

OpenStudy (anonymous):

good friend!!!

OpenStudy (anonymous):

I like the second method.. and here it goes \[f(x)=\frac{x-13}{x^2-25}=(x-13)(x^2-25)^{-1}\\ f'(x)=(x-13)\left[(-1)(x^2-25)^{-2}{d\over dx}(x^2-25)\right]+(1)(x^2-25)^{-1}\\ f'(x)={-2x(x-13)\over(x^2-25)^2}+{1\over x^2-25}\]

OpenStudy (anonymous):

you may simplify by taking the common denominator

OpenStudy (anonymous):

\[\frac{ (x-1)(x-25) }{ (x^2-25)^2 }\]

OpenStudy (anonymous):

how to find its stationary point???

OpenStudy (anonymous):

at a stationary point \(f'(x)=0\)

OpenStudy (anonymous):

at x=1, x=25 and x =5 is that right???

OpenStudy (anonymous):

there seem to be only two values, not 3

OpenStudy (anonymous):

which one is incorrect???

OpenStudy (anonymous):

x=5 are incorrect right

OpenStudy (unklerhaukus):

right

OpenStudy (anonymous):

thank u all